6.if \\(f(x)=2x-4)\\ and \\(g(x)= x^{2} + 3)\\, find f(g(x))
a.\\(-2x^{2} +2)\\
b.\\(-2x^{2} -2)\\
c.\\(2x^{2} -2)\\
d.\\(2x^{2} +2)\\

7.\\(f^{-1}(x)= x - 8)\\, then f(x)=?
a.x/-8
b.1/(x-8)
c.X+8
d.\\( (x-8)^{-1})\\
1
Expert's answer
2019-09-11T13:26:50-0400

"6.\\f(x)=2x-4"

"g(x)=x^2+3"

Now,

"f(g(x))=f(x^2+3)=2(x^2+3)-4=2x^2+6-4"

"f(g(x))=2x^2+2"

Hence option (d) is correct


"7.\\f^{-1}(x)=x-8\\Let\ f^{-1}(x)=y\\So,\\f(y)=x\ \ \ \ \ \ \ \ ..... \ \ \ (1)"

"y=x-8\\x=y+8\\f(y)=y+8\\
f(x)=x+8"

So,

option (c) is correct